OpenStudy (anonymous):

Determine whether the lines are parallel, intersect, or coincide 3x+2y=5 and 6x-4y=8

OpenStudy (texaschic101):

3x + 2y = 5 2y = -3x + 5 y = -3/2x + 5/2 -- slope is -3/2 6x - 4y = 8 -- divide by 2 3x - 2y = 4 -2y = 3x + 4 y = -3/2x - 2 -- slope is -3/2 same slope different y intercepts means the lines are parallel.